Joanne makes a mixture of dried fruits by mixing dried apples costing $6.00/kg and dried apricots costing $8.00/kg. How many kilograms of each are needed to make 20 kg of a mixture worth $7.20/kg?
Answer by Paul(988)

EquatiON:
6x+8(20-x)=20(7.20)
6x+160-8x=144
-2x=-16
x=8

20-8=12

Hence, for apple costing %6 about 8kg of solution; and for aplricots costing $8 about 12 kg of solution is there.
Paul.